The price of AAPL today is $300 per share. AAPL has committed with certainty

to make a $5 dividend payment in two years. The c.c. risk-free interest rate is zero

percent. Assume there is no arbitrage. What is the one-year forward price of AAPL?
Answer #1

One year forward price=(Current spot price-Present Value of Dividends)*(1+risk free rate)^1=(300-5)=295